# FoREST: Frame of Reference Evaluation in Spatial Reasoning Tasks

### v2

@@ Appendix F VISOR-uncond Score
VISORuncond provides the overall spatial relation score, including images with object generation errors. Since it is less focused on evaluating spatial interpretation than VISORcond, which assesses explicitly the text-to-image model’s spatial reasoning, we report VISORuncond results here in the Table 10 rather than in the main paper. The results are similar to the pattern observed in VISORuncond that the based models(SD-1.5 and SD-2.1) perform better in the relative frame of reference, while the layout-to image models, i.e. GLIGEN, are better in the intrinsic frame of reference.
@@
Table 10: VISORuncond score on the I-A-Split and I-C-Split where $I$ refer to the Cow Case and Car Case where relatum has intrinsic directions, and $R$ refer to the Box Case and Pen case where relatum lacks intrinsic directions, $avg$ is mirco-average of $I$ and $R$. cond and uncond are explained in Section 5.1.
Appendix G Analyze the improvements in SG-prompting for T2I.
@@
Table 11: Layout accuracy where spatial relations are left or right relative to the camera. Layout is evaluated for all generated layouts in I-C split while Layoutcond uses the same testing examples as VISORcond.
@@
To further explain improvements of SG-prompting in T2I task, we assess the generated bounding boxes in the I-C split for left and right relations relative to the camera since these can be evaluated using only bounding boxes without depth information. As seen is Table 11, our SG prompting improved Llama3-70B’s by $3.48\%$, while Llama3-8B saw a slight decrease of $0.22\%$. This evaluation was conducted on all generated layouts from the I-C split, which differs from the evaluation subset of images used for VISORcond in Table 3. We report the layoutcond score for a consistent evaluation in the same table. Layoutcond shows that Llama3-8B improves within the same evaluation subset with VISORcond. Overall, by incorporating FoR information through SG layout diffusion, Llama3 generates better spatial configurations, enhancing image generation performance.
Appendix H Frame of Reference Identification
@@
We evaluate the LLMs’ performance in recognizing the FoR classes from given spatial expressions. The LLMs receive spatial expression, denoted as $T$, and output one FoR class, $FoR$, from the valid set of FoR classes, $For\in$external relative, external intrinsic, internal intrinsic, internal relative}. All in-context learning examples are in the Appendix K.
H.1 Experimental Setting
@@
Zero-shot model. We follow the regular setting of zero-shot prompting. We only provide instruction to LLM with spatial context. The instruction prompt briefly explains each class of the FoR and candidate answers for the LLM. We called the LLM with the instruction prompt and $T$ to find $F$.
@@
Few-shot model. We manually craft four spatial expressions for each FoR class. To avoid creating bias, each spatial expression is ensured to fit in only one FoR class. These expressions serve as examples of our few-shotsetting. We provide these examples in addition to the instruction as a part of the prompt, followed by $T$ and query $F$ from the LLM.
@@
Chain-of-Thought (CoT) model. To create CoT (Wei et al., 2023) examples, we modify the prompt to require reasoning before answering. Then, we manually crafted reasoning explanations with the necessary information for each example used in few-shot. Finally, we call the LLMs, adding modified instructions to updated examples, followed by $T$ and query $F$.
Spatial-Guided Prompting (SG) model. We hypothesize that the general spatial relation types defined in Section 2 can provide meaningful information for recognizing FoR classes. For instance, a topological relation, such as “inside," is intuitively associated with an internal FoR. Therefore, we propose Spatial-Guided Prompting to direct the model in identifying the type of relations before querying $F$. We revise the prompting instruction to guide the model in considering these three aspects. Then, we manually explain these three aspects. We specify the relation’s origin from the context for direction relations, such as "the left direction is relative to the observer." We hypothesize that this information helps the model distinguish between intrinsic and relative FoR. Next, we specify whether the locatum is inside or outside the relatum for topological relations. This information should help distinguish between internal and external FoR classes. Lastly, we provide the potential quantitative distance, e.g., far. This quantitative distance further encourages identifying the correct topological and directional relations. Eventually, we insert these new explanations in examples and call the model with the updated instructions followed by $T$ to query $F$.
H.2 Evaluation Metrics
@@
Table 12: The comparison between CoT and SG prompting in C-split separated by inherently clear / required template to be clear.
@@
Model ER-C-Split EC-Split IC-Split IR-C-Split Avg.
@@
Gemma2-9B (4-shot) $59.58$ $55.89$(${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\downarrow 38.34}$) $72.61$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 36.63}$) $74.22$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 20.31}$) $54.44$(${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\downarrow 3.23}$) $64.29$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 3.84}$)
Gemma2-9B (CoT) $60.49$ $60.49$(${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\downarrow 33.74}$) $60.54$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 24.57}$) $87.50$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 33.59}$) $54.03$(${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\downarrow 3.63}$) $65.64$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 5.20}$)
Gemma2-9B (SG)(Our) $72.67$ $65.87$(${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\downarrow 28.37}$) $65.54$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 29.57}$) $53.12$(${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\downarrow 0.78}$) $\mathbf{95.97}$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 38.31}$) $70.13$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 9.68}$)
@@
llama3-8B (4-shot) $60.14$ $47.77$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 15.58}$) $54.35$(${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\downarrow 35.76}$) $100.00$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 24.22}$) $41.13$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 41.13}$) $60.81$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 11.29}$)
llama3-8B (CoT) $61.32$ $61.06$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 28.86}$) $97.28$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 7.17}$) $100.00$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 24.22}$) $36.29$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 36.29}$) $73.66$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 24.14}$)
llama3-8B (SG) (Our) $62.95$ $63.29$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 31.09}$) $94.57$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 4.46}$) $100.00$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 24.22}$) $43.55$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 43.55}$) $75.35$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 25.83}$)
@@
llama3-70B (4-shot) $78.47$ $81.81$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 7.72}$) $64.89$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 55.33}$) $100.00$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 7.81}$) $75.81$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 7.26}$) $80.63$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 19.53}$)
llama3-70B (CoT) $69.11$ $72.05$(${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\downarrow 2.03}$) $97.07$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 87.50}$) $100.00$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 7.81}$) $79.44$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 10.89}$) $87.14$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 26.04}$)
llama3-70B (SG) (Our) $76.50$ $78.21$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 4.12}$) $97.61$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 88.04}$) $100.00$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 7.81}$) $72.18$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 3.63}$) $87.00$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 25.90}$)
@@
Qwen2-7B (4-shot) $61.12$ $50.52$(${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\downarrow 29.45}$) $65.76$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 6.52}$) $93.75$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 16.41}$) $56.05$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 15.32}$) $66.52$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 2.20}$)
Qwen2-7B (CoT) $72.12$ $70.81$(${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\downarrow 9.16}$) $63.80$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 4.57}$) $99.22$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 21.88}$) $51.61$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 10.89}$) $71.36$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 7.04}$)
Qwen2-7B (SG) $70.61$ $68.00$(${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\downarrow 11.98}$) $71.20$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 11.96}$) $88.28$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 10.94}$) $57.26$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 16.53}$) $71.18$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 6.86}$)
@@
Qwen2-72B (4-shot) $79.12$ $78.73$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 16.03}$) $99.35$(${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\downarrow 0.65}$) $87.50$(${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\downarrow 12.50}$) $87.10$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 47.98}$) $88.17$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 12.72}$)
Qwen2-72B (CoT) $88.54$ $88.87$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 26.18}$) $89.57$(${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\downarrow 10.43}$) $93.75$(${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\downarrow 6.25}$) $83.47$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 44.35}$) $88.91$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 13.46}$)
Qwen2-72B (SG) $90.51$ $90.18$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 27.49}$) $93.26$(${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\downarrow 6.74}$) $98.44$(${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\downarrow 1.56}$) $85.08$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 45.97}$) $91.74$(${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 16.29}$)

### v3

@@ Appendix I VISOR-uncond Results
As mentioned in Section 5.1, VISORuncond provides the overall spatial relation score, including images with object generation errors. Because it is less focused on evaluating spatial interpretation than VISORcond, which explicitly assesses a text-to-image model’s spatial reasoning, we report VISORuncond results here in Table 16 rather than in the main paper. The results follow a similar pattern with the VISORcond metric, that is, base models (SD-1.5 and SD-2.1) perform better under the relative frame of reference, while layout-to-image models (e.g., GLIGEN) perform better under the intrinsic frame of reference.
@@
Table 16: VISORuncond score on the A-Split and C-Split where II refers to the Cow Case and Car Case, where relatum has intrinsic directions, and $R$ refers to the Box Case and Pen case, where relatum lacks intrinsic directions, avgavg is the micro-average of II and $R$.
Appendix J Analysis of SG-prompting in T2I
@@
Table 17: Layout accuracy where spatial relations are left or right relative to the camera. Layout is evaluated for all generated layouts in C-split while Layoutcond uses the same testing examples as VISORcond.
@@
To further explain the improvements of SG-prompting in the T2I task, we assess generated bounding boxes in the C-split for left–right relations relative to the camera, since these can be evaluated using bounding boxes alone without depth information. As shown in Table 17, SG-prompting improved Llama3-70B’s performance by 3.48%, while Llama3-8B saw a slight decrease of 0.22%. This evaluation uses all generated layouts from the C-split, differing from the image subset used for VISORcond in Table 4. For consistency, we also report the layoutcond score in Table 17, which shows that Llama3-8B improves within the same evaluation subset as VISORcond. Overall, incorporating FoR information through SG layout diffusion enables Llama3 to generate better spatial configurations, thereby enhancing image generation performance.
Appendix K Frame of Reference Identification Task
@@
We evaluate the LLMs’ performance in recognizing the FoR classes from given spatial expressions. Each model receives a spatial expression $T$ and outputs one FoR class $FoR$ from the valid set of FoR classes, $For\in$ {external relative, external intrinsic, internal intrinsic, internal relative}. All in-context learning examples are in the Appendix L.
K.1 Experimental setting
@@
Zero-shot model. We follow the regular setting of zero-shot prompting. We called the LLM with the instruction prompt and $T$ to find the corresponding FoR, $F$, of given $T$.
@@
Few-shot model. We manually craft four spatial expressions for each FoR class. To avoid creating bias, each spatial expression is ensured to fit in only one FoR class. We provide these examples in addition to the instruction as a part of the prompt, followed by $T$ and query $F$ from the LLM.
@@
Chain-of-Thought (CoT) model. To create CoT (Wei et al., 2022) examples, we modify the prompt to require reasoning before answering. Then, we manually crafted reasoning explanations for each example used in the few-shot. Finally, we call the LLMs, adding modified instructions to updated examples, followed by $T$ and query $F$.
Spatial-Guided Prompting (SG) model. This follows the same SG setting described in Section 4.3. We prompt the LLM to extract spatial information from the given $T$, expecting it to return the FoR as part of the SG response. We then extract this FoR to obtain $F$. Unlike QA and T2I tasks, this FoR identification task does not invoke additional CoT for downstream reasoning.
K.2 Evaluation metrics
@@
Table 18: Comparison of CoT and SG prompting in the C-split, separated into inherently clear contexts (without FoR ambiguity) and template-dependent contexts (requiring extra information to resolve ambiguity).
@@
Model ER-Split EI-Split II-Split IR-Split Avg.
@@
Gemma2-9B (4-shot) $59.58$ 55.8955.89($\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\downarrow 38.34$}) 72.6172.61($\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 36.63$}) 74.2274.22($\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 20.31$}) 54.4454.44($\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\downarrow 3.23$}) 64.2964.29($\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 3.84$})
Gemma2-9B (CoT) $60.49$ 60.4960.49($\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\downarrow 33.74$}) 60.5460.54($\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 24.57$}) 87.5087.50($\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 33.59$}) 54.0354.03($\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\downarrow 3.63$}) 65.6465.64($\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 5.20$})
Gemma2-9B (SG)(Our) $72.67$ 65.8765.87($\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\downarrow 28.37$}) 65.5465.54($\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 29.57$}) 53.1253.12($\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\downarrow 0.78$}) $\mathbf{95.97}$($\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 38.31$}) 70.1370.13($\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 9.68$})
@@
llama3-8B (4-shot) $60.14$ 47.7747.77($\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 15.58$}) 54.3554.35($\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\downarrow 35.76$}) 100.00100.00($\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 24.22$}) 41.1341.13($\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 41.13$}) 60.8160.81($\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 11.29$})
llama3-8B (CoT) $61.32$ 61.0661.06($\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 28.86$}) 97.2897.28($\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 7.17$}) 100.00100.00($\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 24.22$}) 36.2936.29($\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 36.29$}) 73.6673.66($\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 24.14$})
llama3-8B (SG) (Our) $62.95$ 63.2963.29($\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 31.09$}) 94.5794.57($\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 4.46$}) 100.00100.00($\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 24.22$}) 43.5543.55($\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 43.55$}) 75.3575.35($\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 25.83$})
@@
llama3-70B (4-shot) $78.47$ 81.8181.81($\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 7.72$}) 64.8964.89($\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 55.33$}) 100.00100.00($\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 7.81$}) 75.8175.81($\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 7.26$}) 80.6380.63($\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 19.53$})
llama3-70B (CoT) $69.11$ 72.0572.05($\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\downarrow 2.03$}) 97.0797.07($\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 87.50$}) 100.00100.00($\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 7.81$}) 79.4479.44($\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 10.89$}) 87.1487.14($\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 26.04$})
llama3-70B (SG) (Our) $76.50$ 78.2178.21($\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 4.12$}) 97.6197.61($\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 88.04$}) 100.00100.00($\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 7.81$}) 72.1872.18($\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 3.63$}) 87.0087.00($\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 25.90$})
@@
Qwen2-7B (4-shot) $61.12$ 50.5250.52($\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\downarrow 29.45$}) 65.7665.76($\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 6.52$}) 93.7593.75($\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 16.41$}) 56.0556.05($\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 15.32$}) 66.5266.52($\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 2.20$})
Qwen2-7B (CoT) $72.12$ 70.8170.81($\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\downarrow 9.16$}) 63.8063.80($\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 4.57$}) 99.2299.22($\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 21.88$}) 51.6151.61($\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 10.89$}) 71.3671.36($\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 7.04$})
Qwen2-7B (SG) $70.61$ 68.0068.00($\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\downarrow 11.98$}) 71.2071.20($\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 11.96$}) 88.2888.28($\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 10.94$}) 57.2657.26($\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 16.53$}) 71.1871.18($\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 6.86$})
@@
Qwen2-72B (4-shot) $79.12$ 78.7378.73($\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 16.03$}) 99.3599.35($\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\downarrow 0.65$}) 87.5087.50($\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\downarrow 12.50$}) 87.1087.10($\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 47.98$}) 88.1788.17($\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 12.72$})
Qwen2-72B (CoT) $88.54$ 88.8788.87($\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 26.18$}) 89.5789.57($\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\downarrow 10.43$}) 93.7593.75($\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\downarrow 6.25$}) 83.4783.47($\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 44.35$}) 88.9188.91($\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 13.46$})
Qwen2-72B (SG) $90.51$ 90.1890.18($\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 27.49$}) 93.2693.26($\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\downarrow 6.74$}) 98.4498.44($\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\downarrow 1.56$}) 85.0885.08($\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 45.97$}) 91.7491.74($\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\uparrow 16.29$})